this is a differential equation question please solve them with steps and details also, make sure it is correct Find the following functions pleas make your font clear

(b) C-H9_2 2s +1 (c) L-{ $2 + 6s + 25 }
Answer #1

e 45 S-2 b) { u(t-1) € 214-4). [:<{s-a} = cat] Ult-4)e?t-8 c) L L = 1 25 + 1 s? +65 +25 } 2 S +1 5+65+9+16 tie} = . ב 5 }